# Curriculum-style Data Augmentation for LLM-based Metaphor Detection

### v1

@@ Abstract
Kaidi Jiathanks: Equal contribution, Yanxia Wu*  and Rongsheng Li*thanks: Corresponding author College of Computer Science and Technology, Harbin Engineering University, Harbin 150001, China dasheng@hrbeu.edu.cn
@@ 1 Introduction
Figure 1: Structures of Our Method
@@ 5.1 Datasets
We fine-tune the model on the train dataset using CDA and perform zero-shot testing on the test dataset. The statistical information of the dataset is shown in Table 3.
5.1.1 Train Datasets
VUA Verb (Steen et al., 2010): The dataset is a verb subset of the largest metaphor detection dataset, VUA ALL (Steen et al., 2010), collected by VUA from the BNCBaby corpus. Due to the extremely imbalanced data distribution, we selected 6000 instances for training, consisting of 3000 metaphorical instances and 3000 literal instances.
@@ 5.1.2 Test Datasets
For comparison with methods based on LLMs, we follow the approach of Tian et al. (2024) and select 300 instances from each of the two test datasets for testing, including 150 metaphorical instances and 150 literal instances.
@@ 5.6.5 Ablation for Different Augmentation Methods
  w.Direct 86.00 86.09 68.00 72.73
w.Target 83.33 82.76 69.00 73.50
w.Context 84.33 84.28 67.67 72.21

### v2

@@ Abstract
Kaidi Jia1111thanks: Equal contribution, Yanxia Wu11*1 ∗, Ming Liu2222  and Rongsheng Li11*1 ∗thanks: Corresponding author $1$College of Computer Science and Technology, Harbin Engineering University, Harbin 150001, China $2$Deakin University, Australia dasheng@hrbeu.edu.cn
@@ 1 Introduction
Figure 1: Structures of Our Method. The top half shows the process of Curriculum-style Data Augmentation (CDA), and the bottom half shows examples of data augmentation, metaphorical examples in blue boxes and non-metaphorical examples in green boxes.
@@ 5.1 Datasets
We fine-tune the model on the train dataset using CDA and perform zero-shot testing on the test dataset. The statistical information of the dataset is shown in Table 2. The statistical information of the dataset after augmented can be found in the appendix B.
5.1.1 Train Datasets
VUA Verb (Steen et al., 2010): The dataset is a verb subset of the largest metaphor detection dataset, VUA ALL (Steen et al., 2010), collected by VUA from the BNCBaby corpus. The initial VUA Verb dataset comprised a total of 3,251 metaphorical instances and 10,357 non-metaphorical instances. To ensure a balanced distribution of the initial training data, we randomly selected 3,000 metaphorical instances and 3,000 non-metaphorical instances. The random selection was performed using a fixed seed of 42 to ensure reproducibility.
@@ 5.1.2 Test Datasets
For comparison with methods based on LLMs, we directly used the data selected from TSI (Tian et al., 2024) without any modification, with 150 metaphorical instances and 150 non-metaphorical instances extracted from each dataset.
While our experiments were conducted only on the verb datasets, this does not imply that our method is exclusively applicable to verb metaphor detection. We extended our experimental scope to other parts of speech within the VUA ALL dataset, encompassing adjectives, adverbs, and nouns, which can be found in appendix C.
@@ 5.6 Ablation Study
5.6.5 Ablation for CDA
  Dataset Using CDA Acc. F1
MOH-X 85.00 85.98
86.11 86.57
TroFi 65.67 70.99
69.11 73.62
Table 9: Ablation study on CDA. Using CDA refers to fine-tuning using CDA vs. fine-tuning using LoRA only.
In order to explore the overall effect of CDA, we conducted an ablation experiment on CDA. We compared the performance of fine-tuning using CDA versus fine-tuning using LoRA only, as shown in Table 9. On both datasets, the effect of fine-tuning using CDA was much higher than that of fine-tuning using LoRA only, suggesting that our improvement was mainly due to CDA rather than fine-tuning.
@@ 5.6.6 Ablation for Different Augmentation Methods
  w/o Direct 86.00 86.09 68.00 72.73
w/o Target 83.33 82.76 69.00 73.50
w/o Context 84.33 84.28 67.67 72.21
@@ Appendix A Prompts
Appendix B Statistical Information of The Augmented Data
  Iteration #Instance %M %Correct %Correct.M
I 6,000 50.00 68.62 56.01
II 11,416 44.48 87.99 48.56
III 15,324 36.88 88.72 39.92
  Method VUA Adj. VUA Adv. VUA Noun.
Acc. F1 Acc. F1 Acc. F1
  Llama 3.1 8B Instruct 55.84 29.31 55.47 16.01 62.74 34.21
Llama 3.1 8B Instruct (CDA) 76.14 42.46 85.88 35.88 80.64 52.39
Table 18: Zero-shot experiments conducted on the VUA ALL dataset.
Appendix C Zero-shot Transfer on VUA ALL
We extended our experimental scope to other parts of speech within the VUA ALL dataset, encompassing adjectives, adverbs, and nouns. It is important to note that these additional experiments were conducted in a zero-shot manner, meaning that the model was trained exclusively on the verb dataset and then directly tested on datasets of other parts of speech. The experimental results are presented in Table 18. As can be seen, even when trained solely on the verb dataset, the model’s performance on each part-of-speech dataset is significantly enhanced. Thus, our approach is not limited to verb datasets.